Suppose you analyzed the average monthly credit card bill of Visa credit card customers using two randomly selected samples: one with 133 Visa credit card customers (Sample #1) and one with 108 Visa credit card customers (Sample #2). Every credit card customer in Sample #1 had a monthly credit card bill of $1,831. Sample #2, on the other hand, had an average monthly credit card bill of $1,871 with a variance of 1, 09,043(s 12 fro combine these samples into one... Question 2 What would be the combined monthly credit card bill average (in dollars)? Question 3A What is the combined monthly credit card bill median (in dollars)? Question 3BWhat is the shape of the combined monthly credit card bill distribution (in dollars)? Question 4 What would be the combined monthly credit card bill standard deviation (in dollars)?please CARRY ALL DECIMALS!!
Answer #1

ques 2

sample 1

n1=133

xbar 1=1831

s1=0

sample 2

n2=108

xbar 2=1871

s2² = 1309049

cpmbined average = [n1 x1bar + n2 x2bar]/[n1+n2] = (133*1831+108*1871)/(133+108) = 1848.9253

question 3 A

The total number of visa card customers is 133+108=241 cards here.

Thus, the median is (241+1)/2=121th value.

As all 133 visa card customers have bill is 1831,

The median is,$1831

question 3B

since mean>median ,so

the shape of the combined monthly credit card bill is skewed to right

3C

Use

s= + n2-2 to find the standard deviation of combined monthly credit card bill.

s^2 = =(107*1309049+132*(1831-1848.925)^2+107*(1871-1848.925)^2)/(133+108-2) = 586455.2178

so, s = √s^2 = $765.8036